Please help me solve this question. A student answers 80% of the questions on a math exam correctly. If she answers 20 questions correctly,how many questions are on the exam?
Please help!
Thanks,
Adalia

Answer by wilft1(217)

another way of asking this question would be... "20 is 80 percent of what number" ?
80/100 = 20/x
reduce to lowest terms...
4/5 = 20/x
this makes it a little easier to figure out...
divide 20 by 4, then multiply by 5 to get 25
so there were 25 questions on the test